In a sample of 6000 new iPhones, 48 were found to be defective. At this rate, how many defective iPhones would be expected in a sample of 35,000 iPhones?
Found 2 solutions by rfer, ikleyn:
Answer by rfer(16322)

48/6000=x/35000
6000x=1680000
x=1680000/6000
x=280 defective

Answer by ikleyn(52845)

.

See the classic problem of the same type in the lesson
    - Using proportions to estimate the number of fish in a lake
in this site.

When you read this lesson, you will not miss this kind of problems with anything else.